What is the angular speed in rad/s of body whose frequency is 50Hz
1
Expert's answer
2020-12-03T10:25:29-0500

We can find the angular speed of body from the formula:


"\omega=2\pi f=2\cdot\pi \cdot 50\ Hz=314\ \dfrac{rad}{s}."

Answer:

"\omega=314\ \dfrac{rad}{s}."